# von Neumann Orbit Equivalence

### v1

@@ 1. Introduction
Let $\Gamma$ and $\Lambda$ be two countable discrete groups with free probability measure-preserving actions $\Gamma{\,\curvearrowright\,}(X,\mu)$ and $\Lambda{\,\curvearrowright\,}(Y,\nu)$ on standard probability measure spaces $(X,\mu)$ and $(Y,\nu)$, respectively. An orbit equivalence (OE) for the actions is a measurable isomorphism $\theta:X\to Y$ such that $\theta(\Gamma x)=\Lambda\theta(x)$ for almost every $x\in X$. In this case, the two actions are called orbit equivalent. Two groups are said to be orbit equivalent if they admit orbit equivalent actions. Singer [Sin55] showed that OE of $\Gamma{\,\curvearrowright\,}(X,\mu)$ and $\Lambda{\,\curvearrowright\,}(Y,\nu)$ is equivalent to the existence of an isomorphism $L^{\infty}(X)\rtimes\Gamma\cong L^{\infty}(Y)\rtimes\Lambda$ which preserves the Cartan subalgebras $L^{\infty}(X)$ and $L^{\infty}(Y)$. Orbit equivalence theory saw some development in the 1980’s (see [OW80, CFW81, Zim84]), and has been an area of active research over the last two decades (see [Fur11, Gab10]). These advances in part have been stimulated by the success of the deformation/rigidity theory approach to the classification of $\mathrm{II}_{1}$ factors developed by Popa and others (see [Pop08, Vae10, Ioa13]).
The study of orbit equivalence can be motivated also from an entirely different point of view, being a measurable counterpart to quasi-isometry of groups. Gromov [Gro93] introduced measure equivalence (ME) for countable discrete groups as a measurable analogue of quasi-isometry and since then this notion has proven to be an important tool in geometric group theory with connections to ergodic theory and operator algebras. Two infinite countable discrete groups $\Gamma$ and $\Lambda$ are measure equivalent if there is an infinite measure space $(\Omega,m)$ with a measurable, measure-preserving action of $\Gamma\times\Lambda$, so that both the actions $\Gamma{\,\curvearrowright\,}(\Omega,m)$ and $\Lambda{\,\curvearrowright\,}(\Omega,m)$ admit finite-measure fundamental domains $Y,X\in\Omega$:
@@
The space $(\Omega,m)$ is called an ME-coupling of $\Gamma$ with $\Lambda$, and the index of such a coupling is defined as
@@
Notably, measure equivalence was used by Furman in [Fur99a, Fur99b] to prove strong rigidity results for lattices in higher rank simple Lie groups. ME relates back to OE because of the following fact, observed by Zimmer and Furman: for two discrete groups $\Gamma$ and $\Lambda$, admitting free OE actions is equivalent to having an ME-coupling of index 1. Moreover, for OE groups, an ME-coupling can be chosen so that the fundamental domains coincide.
If $\Gamma$ and $\Lambda$ have commuting actions on $(\Omega,m)$ and if $X\subset\Omega$ is a Borel fundamental domain for the action of $\Gamma$, then on the level of function spaces, the characteristic function $1_{X}$ gives a projection in $L^{\infty}(\Omega,m)$ such that the collection $\{1_{\gamma X}\}_{\gamma\in\Gamma}$ forms a partition of unity, i.e., $\sum_{\gamma\in\Gamma}1_{\gamma X}=1$. This notion generalizes quite nicely to the non-commutative setting, and using this, Peterson, Ruth, and the first named author, in [IPR24], defined that a fundamental domain for an action on a von Neumann algebra $\Gamma{\,\curvearrowright^{\sigma}\,}\mathcal{M}$ is a projection $p\in\mathcal{M}$ such that $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(p)=1$, where the convergence is in the strong operator topology. Using this perspective for a fundamental domain they generalized the notion of measure equivalence by simply considering actions on non-commutative spaces.
@@
As mentioned above, OE of groups is equivalent to having an ME-coupling with a common fundamental domain. Motivated by this, we introduce the following definition.
@@
Like ME, vNE is stable under taking the direct product of groups. But neither ME nor vNE is stable under taking free products. For instance, since any two finite groups are ME (and hence vNE), and amenability is preserved under both ME and vNE, one gets that $\mathbb{Z}/2\mathbb{Z}*\mathbb{Z}/2\mathbb{Z}$ (amenable) is neither ME nor vNE to $\mathbb{Z}/3\mathbb{Z}*\mathbb{Z}/2\mathbb{Z}$ (non-amenable). However, as suggested in [MS06, Remark 2.28], and proved in [Gab05, $\mathbf{P}_{\mathrm{ME}}{\bf 6}$], stability under taking free products hold if one requires the additional assumption that groups are ME with a common fundamental domain. In other words, OE is stable under taking free products. This raises a natural question: Is vNOE stable under taking free products? We obtain an affirmative answer to this question.
@@
If $\Gamma_{i},\Lambda_{i},~{}i=1,2$ are countable discrete groups such that $\Gamma_{i}\sim_{\mathrm{vNOE}}\Lambda_{i},~{}i=1,2$, then $\Gamma_{1}*\Gamma_{2}\sim_{\mathrm{vNOE}}\Lambda_{1}*\Lambda_{2}$
Green [Gre90], in her Ph.D. thesis, introduced graph products of groups, another important group theoretical construction. If $\mathcal{G}=(V,E)$ is a simple, non-oriented graph with vertex set $V$ and edge set $E$, then the graph product of a family, $\{\Gamma_{v}\}_{v\in V}$, of groups indexed by $V$ is obtained from the free product $*_{v\in V}\Gamma_{v}$ by adding commutator relations determined by the edge set $E$. Depending on the graph, free products and direct products are special cases of the graph product construction. Adapting the ideas of [Gab05], Horbez and Huang [HH22, Proposition 4.2] proved the stability of OE under taking graph products (see also [Dem22]). We are also able to show that the stability of vNOE under taking graph products. To see more about the study of graph products in the realm of measured group theory, we would like to bring the article [EH24] to the reader’s attention.
@@
Let $\mathcal{G}=(V,E)$ be a simple finite graph. Let $\Gamma$ and $\Lambda$ be two graph products over $\mathcal{G}$, with countable vertex groups $\{\Gamma_{v}\}_{v\in V}$ and $\{\Lambda_{v}\}_{v\in V}$, respectively. If $\Gamma_{v}\sim_{\mathrm{vNOE}}\Lambda_{v}$ for every $v\in V$, then $\Gamma\sim_{\mathrm{vNOE}}\Lambda$.
In attempting to prove the above theorems, if one tries to adapt the techniques from one of [Gab05, HH22, Dem22], an immediate problem is presented by the lack of “point perspective” in the theory of von Neumann (orbit) equivalence. The lack of any natural non-commutative analogue of the notion of OE/ME cocycles, or that of measured equivalence relation can be considered as a few problems presented by the lack of point perspective. This often leads one to consider genuinely new techniques and different alternatives (see e.g., [IPR24, Ish24, Bat23a, Bat23b]). To overcome this obstruction, we introduce the notion of von Neumann orbit equivalence for tracial von Neumann algebras that is “compatible” with vNOE of groups (see Theorem 3.9), and prove the analogues of Theorems 1.3 and 1.4 at the level of tracial von Neumann algebras.
The notion of von Neumann equivalence admits a generalization in the setting of finite von Neumann algebras [IPR24, Section 8], and relates to vNE for groups as follows: $\Gamma\sim_{\mathrm{vNE}}\Lambda$ if and only if $L\Gamma\sim_{\mathrm{vNE}}L\Lambda$ [IPR24, Theorem 1.5]. In parallel to this, one might attempt to define two tracial von Neumann algebras to be vNOE if they are vNE and admit a “common” fundamental domain, and identify a correct meaning of “common”. However, we take a slightly different approach, and motivated by the recently defined notion of measure equivalence of finite von Neumann algebras by Berendschot and Vaes in [BV22], we introduce the following definition.
@@
Let $(A,\tau_{A})$ and $(B,\tau_{B})$ be tracial von Neumann algebras. We say that $A$ and $B$ are von Neumann orbit equivalent, denoted $A\sim_{\mathrm{vNOE}}B$, if there exists a tracial von Neumann algebra $(Q,\tau_{Q})$, a Hilbert $A\,\overline{\otimes}\,Q-B$-bimodule $\mathcal{H}$, and a vector $\xi\in\mathcal{H}$ such that
@@
    $\langle(a\otimes x)\xi,\xi\rangle=\tau_{A}(a)\tau_{Q}(x)$, and $\langle y\xi b,\xi\rangle=\tau_{Q}(y)\tau_{B}(b)$ for every $a\in A,x,y\in Q,$ and $b\in B$.
@@
We prove in Proposition 3.4 that vNOE is indeed an equivalence relation. We should remark that, in the above definition, $\mathcal{H}$ can also be considered as an $A-B\,\overline{\otimes}\,Q^{\mathrm{op}}$-bimodule satisfying conditions analogous to the two mentioned in the definition. This essentially is the reason for the symmetry of vNOE, even though the formal definition seems asymmetric. To prove transitivity, inspired by [BV22, Lemma 5.11], we establish an equivalent characterization of vNOE in Theorem 3.1, and show in Theorem 3.9 that for countable discrete groups $\Gamma$ and $\Lambda$, $\Gamma\sim_{\mathrm{vNOE}}\Lambda$ if and only if $L\Gamma\sim_{\mathrm{vNOE}}L\Lambda$. Since $L(\Gamma*\Lambda)\cong L\Gamma*L\Lambda$, Theorem 1.3 will follow from the following, which we prove in Section 3.
@@
Let $A_{i},B_{i},~{}i=1,2$ be tracial von Neumann algebras such that $A_{i}\sim_{\mathrm{vNOE}}B_{i},~{}i=1,2$. Then, $A_{1}*A_{2}\sim_{\mathrm{vNOE}}B_{1}*B_{2}$.
Similar to free products, one also has that the group von Neumann algebra of a graph product of groups is isomorphic to the (von Neumann algebra) graph product of the group von Neumann algebras, and hence Theorem 1.4 follows from the following theorem, proved in Section 4.
@@
Let $\mathcal{G}=(V,E)$ be a simple finite graph. Let $A$ and $B$ be two graph products over $\mathcal{G}$, with tracial vertex von Neumann algebras $\{A_{v}\}_{v\in V}$ and $\{B_{v}\}_{v\in V}$, respectively. If $A_{v}\sim_{\mathrm{vNOE}}B_{v}$ for every $v\in V$, then $A\sim_{\mathrm{vNOE}}B$.
In Proposition 3.12, we show that vNOE tracial von Neumann algebras are indeed vNE in the sense of [IPR24]. We should remark that vNE does not imply vNOE in general.
@@ Acknowledgements.
The authors would like to express deep gratitude to Ben Hayes for his valuable insights, continuous support, and constant encouragement throughout this project. This work has benefitted profoundly from discussions with Ben Hayes on numerous occasions and from him encouraging us to look into [BV22]. The project began with the authors’ meeting at the “NCGOA Spring Institute 2023” held at the Vanderbilt University, and we thank the organizers of the conference for the opportunity. A.Wu would also like to acknowledge the support from the NSF CAREER award DMS ##\##214473.
@@ 2.1. Tracial von Neumann algebras and the standard form
2.2. Graph product
Let $\mathcal{G}=(V,E)$ be a simple graph with the vertex set $V$ and the edge set $E\subseteq V\times V\setminus\{(v,v):v\in V\}$. We assume that the graph $\mathcal{G}$ is non-oriented, i.e., $(v,w)\in E$ if and only if $(w,v)\in E$. For a vertex $v\in V$, we let $\mathsf{lk}(v)$ denote the set of all vertices that are connected to $v$ by an edge, i.e., $\mathsf{lk}(v)=\{w\in V:(v,w)\in E\}$; and we let $\mathsf{st}(v)=\mathsf{lk}(v)\cup\{v\}$. If $U\subseteq V$, then the full subgraph of $\mathcal{G}$ with the vertex set $U$ is the graph with $U$ as the vertex set and $v,w\in U$ are connected by an edge if and only if $(v,w)\in E$. Abusing the notation slightly, we will use $\mathsf{lk}(v)$ and $\mathsf{st}(v)$ to denote the full subgraph of $\mathcal{G}$ with vertex sets $\mathsf{lk}(v)$ and $\mathsf{st}(v)$, respectively.
A word $v_{1}v_{2}\cdots v_{n}$ of vertices in $V$ is called reduced if it satisfies the following property: if there exist $k<l$ such that $v_{k}=v_{l}$, then there is some $k<j<l$ such that $(v_{k},v_{j})\notin E$. Let $\mathcal{G}=(V,E)$ be a simple graph, $(A,\tau)$ be a tracial von Neumann algebra, and $\{(A_{v},\tau_{v}):v\in V\}$ be a family of tracial von Neumann subalgebras of $(A,\tau)$ such that $\tau|_{A_{v}}=\tau_{v}$ for all $v\in V$. We say that the family $\{(A_{v},\tau_{v}):v\in V\}$ is $\mathcal{G}$-independent if the following property holds: if $v_{1}\cdots v_{n}$ is a reduced word and $a_{1},\ldots,a_{n}\in A$ are such that $a_{i}\in A_{v_{i}}$ and $\tau(a_{i})=0$, then $\tau(a_{1}\cdots a_{n})=0$. On the other hand, given a simple graph $\mathcal{G}=(V,E)$ and a family of tracial von Neumann algebras $\{(A_{v},\tau_{v}):v\in V\}$, there is a unique, up to isomorphism, tracial von Neumann algebra $(A,\tau)$, called the graph product over $\mathcal{G}$ of the family $\{(A_{v},\tau_{v}):v\in V\}$, and trace-preserving inclusions $\varphi_{v}:A_{v}\hookrightarrow A$ such that the family $\{\varphi_{v}(A_{v}):v\in V\}$ is $\mathcal{G}$-independent and generates $A$ as a von Neumann algebra (see [Mło04, CF17]). We denote the graph product $(A,\tau)$ of the family $\{(A_{v},\tau_{v}):v\in V\}$ by
$(A,\tau)=\bigstar_{v\in V}(A_{v},\tau_{v}).$
Given a simple graph $\mathcal{G}=(V,E)$, a family of tracial von Neumann algebras $\{(A_{v},\tau_{v}):v\in V\}$, and a vertex $v\in V$, we let
$A_{\mathsf{st}(v)}=\bigstar_{w\in\mathsf{st}(v)}A_{w},\quad A_{V^{\prime}}=\bigstar_{w\in V\setminus\{v\}}A_{w},~{}~{}\text{ and }~{}~{}A_{\mathsf{lk}(v)}=\bigstar_{w\in\mathsf{lk}(v)}A_{w}.$
Then, by the unscrewing technique of Caspers and Fima ([CF17, Theorem 3.26]), there exists a unique trace-preserving *-isomorphism $\Phi:A_{\mathsf{st}(v)}*_{A_{\mathsf{lk}(v)}}A_{V^{\prime}}\to\bigstar_{v\in V}A_{v}$ such that $\Phi|_{\mathsf{st}(v)}$ and $\Phi|_{A_{V^{\prime}}}$ are the canonical inclusions $A_{\mathsf{st}(v)}\subset A$ and $A_{V^{\prime}}\subset A$, respectively. We remark that one can also view $A$ as $(A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)})*_{A_{\mathsf{lk}(v)}}A_{V^{\prime}}$, a viewpoint that will be useful in proving Theorem 4.1.
Remark 2.1.

### v2

@@ 1. Introduction
Let $\Gamma$ and $\Lambda$ be two countable discrete groups with free probability measure-preserving actions $\Gamma{\,\curvearrowright\,}(X,\mu)$ and $\Lambda{\,\curvearrowright\,}(Y,\nu)$ on standard probability measure spaces $(X,\mu)$ and $(Y,\nu)$, respectively. An orbit equivalence (OE) for the actions is a measurable isomorphism $\theta:X\to Y$ such that $\theta(\Gamma x)=\Lambda\theta(x)$ for almost every $x\in X$. In this case, the two actions are called orbit equivalent. Two groups are said to be orbit equivalent if they admit orbit equivalent actions. Singer [Sin55] showed that for two free probability measure preserving actions $\Gamma{\,\curvearrowright\,}(X,\mu)$ and $\Lambda{\,\curvearrowright\,}(Y,\nu)$, being orbit equivalent is equivalent to the existence of an isomorphism $L^{\infty}(X)\rtimes\Gamma\cong L^{\infty}(Y)\rtimes\Lambda$ which preserves the Cartan subalgebras $L^{\infty}(X)$ and $L^{\infty}(Y)$. Orbit equivalence theory saw some development in the 1980s (see [OW80, CFW81, Zim84]), and has been an area of active research over the last two decades (see [Fur11, Gab10]). These advances in part have been stimulated by the success of the deformation/rigidity theory approach to the classification of $\mathrm{II}_{1}$ factors developed by Popa and others (see [Pop08, Vae10, Ioa13]).
The study of orbit equivalence can be motivated also from an entirely different point of view, being a measurable counterpart to quasi-isometry of groups. Gromov [Gro93] introduced measure equivalence (ME) for countable discrete groups as a measurable analogue of quasi-isometry and since then this notion has proven to be an important tool in geometric group theory with connections to ergodic theory and operator algebras. Two infinite countable discrete groups $\Gamma$ and $\Lambda$ are measure equivalent if there is an infinite measure space $(\Omega,m)$ with commuting, measure-preserving actions $\Gamma{\,\curvearrowright\,}(\Omega,m)$ and $\Lambda{\,\curvearrowright\,}(\Omega,m)$, so that both the actions admit finite-measure fundamental domains $Y,X\in\Omega$, that is, $m(Y),m(X)<\infty$ and
@@
The space $(\Omega,m)$ is called an ME-coupling between $\Gamma$ and $\Lambda$, and the index of such a coupling is
@@
Notably, measure equivalence was used by Furman in [Fur99a, Fur99b] to prove strong rigidity results for lattices in higher rank simple Lie groups. ME relates back to OE because of the following fact, observed by Zimmer and Furman: for two discrete groups $\Gamma$ and $\Lambda$, admitting free OE actions is equivalent to having an ME-coupling of index 1. Moreover, for OE groups, an ME-coupling can be chosen so that the fundamental domains coincide [Fur99b, Theorem 3.3].
If $X\subset\Omega$ is a Borel fundamental domain for the action $\Gamma{\,\curvearrowright\,}(\Omega,m)$, then on the level of function spaces, the characteristic function $1_{X}$ gives a projection in $L^{\infty}(\Omega,m)$ such that the collection $\{1_{\gamma X}\}_{\gamma\in\Gamma}$ forms a partition of unity, i.e., $\sum_{\gamma\in\Gamma}1_{\gamma X}=1$. This notion generalizes quite nicely to the non-commutative setting, and using this, Peterson, Ruth, and the first named author, in [IPR24], defined that a fundamental domain for an action on a von Neumann algebra $\Gamma{\,\curvearrowright^{\sigma}\,}\mathcal{M}$ is a projection $p\in\mathcal{M}$ such that $\sum_{\gamma\in\Gamma}\sigma_{\gamma}(p)=1$, where the convergence is in the strong operator topology. Using this perspective for a fundamental domain they generalized the notion of measure equivalence by considering actions on non-commutative spaces.
@@
Like ME, vNE is stable under taking the direct product of groups. But neither ME nor vNE is stable under taking free products. For instance, since any two finite groups are ME (and hence vNE), and amenability is preserved under both ME and vNE, one gets that $\mathbb{Z}/2\mathbb{Z}*\mathbb{Z}/2\mathbb{Z}$ (amenable) is neither ME nor vNE to $\mathbb{Z}/3\mathbb{Z}*\mathbb{Z}/2\mathbb{Z}$ (non-amenable). However, as suggested in [MS06, Remark 2.28], and proved in [Gab05, $\mathbf{P}_{\mathrm{ME}}{\bf 6}$], stability under taking free products hold if one requires the additional assumption that groups are ME with a common fundamental domain. In other words, OE is stable under taking free products. This raises a natural question: Is vNE, with common fundamental domain, stable under taking free products? We obtain an affirmative answer to this question and introduce the following definition.
@@
If $\Gamma_{i},\Lambda_{i},~{}i=1,2$ are countable discrete groups such that $\Gamma_{i}\sim_{\mathrm{vNOE}}\Lambda_{i},~{}i=1,2$, then $\Gamma_{1}*\Gamma_{2}\sim_{\mathrm{vNOE}}\Lambda_{1}*\Lambda_{2}$.
Remark 1.4.
If $(\mathcal{M},\mathrm{Tr})$ is a von Neumann coupling between $\Gamma$ and $\Lambda$ with $\Gamma$ and $\Lambda$ fundamentals domains $q$ and $p$, respectively, then the index of such a coupling is defined as
$[\Gamma:\Lambda]_{\mathcal{M}}:=\frac{\mathrm{Tr}(p)}{\mathrm{Tr}(q)}$
We suspect that the notion of vNE with coupling index 1 should be equivalent to the notion of vNE with common fundamental domain. However, we are unable to prove it at this point and leave it as an open problem.
Green [Gre90], in her Ph.D. thesis, introduced graph products of groups, another important group theoretical construction. If $\mathcal{G}=(V,E)$ is a simple, non-oriented graph with vertex set $V$ and edge set $E$, then the graph product of a family, $\{\Gamma_{v}\}_{v\in V}$, of groups indexed by $V$ is obtained from the free product $*_{v\in V}\Gamma_{v}$ by adding commutator relations determined by the edge set $E$. Depending on the graph, free products and direct products are special cases of the graph product construction. Adapting the ideas of [Gab05], Horbez and Huang [HH22, Proposition 4.2] proved the stability of OE under taking graph products over finite simple graphs (see also [Dem22]). To further explore the study of graph products within the context of measured group theory, we would like to draw the reader’s attention to the article [EH24]. In the current article, we also prove the stability of vNOE under taking graph products.
@@
Let $\mathcal{G}=(V,E)$ be a simple graph with at most countably infinite vertices. Let $\Gamma$ and $\Lambda$ be two graph products over $\mathcal{G}$, with countable vertex groups $\{\Gamma_{v}\}_{v\in V}$ and $\{\Lambda_{v}\}_{v\in V}$, respectively. If $\Gamma_{v}\sim_{\mathrm{vNOE}}\Lambda_{v}$ for every $v\in V$, then $\Gamma\sim_{\mathrm{vNOE}}\Lambda$.
In attempting to prove the above theorems, if one tries to adapt the techniques from one of [Gab05, HH22, Dem22], an immediate problem is presented by the lack of “point perspective” in the theory of von Neumann (orbit) equivalence. The lack of any natural non-commutative analogue of the notion of OE/ME cocycles, or that of measured equivalence relation can be considered as a few problems presented by the lack of point perspective. This often leads one to consider genuinely new techniques and different alternatives (see e.g., [IPR24, Ish24, Bat23a, Bat23b]). To overcome this obstruction, we introduce the notion of von Neumann orbit equivalence for tracial von Neumann algebras that is “compatible” with vNOE of groups (see Definition 1.6 and Theorem 3.10), and prove the analogues of Theorems 1.3 and 1.5 at the level of tracial von Neumann algebras.
The notion of von Neumann equivalence admits a generalization in the setting of finite von Neumann algebras [IPR24, Section 8], and relates to vNE for groups as follows: $\Gamma\sim_{\mathrm{vNE}}\Lambda$ if and only if $L\Gamma\sim_{\mathrm{vNE}}L\Lambda$ [IPR24, Theorem 1.5]. In parallel to this, one might attempt to define two tracial von Neumann algebras to be vNOE if they are vNE and admit a “common” fundamental domain, and identify a correct meaning of “common”. However, we take a slightly different approach, and motivated by the recently defined notion of measure equivalence of finite von Neumann algebras by Berendschot and Vaes in [BV], we introduce the following definition.
@@
Let $(A,\tau_{A})$ and $(B,\tau_{B})$ be tracial von Neumann algebras. We say that $(A,\tau_{A})$ and $(B,\tau_{B})$ are von Neumann orbit equivalent, denoted $(A,\tau_{A})\sim_{\mathrm{vNOE}}(B,\tau_{B})$, if there exists a tracial von Neumann algebra $(Q,\tau_{Q})$, a Hilbert $A\,\overline{\otimes}\,Q-B$-bimodule $\mathcal{H}$, and a vector $\xi\in\mathcal{H}$ such that
@@
    $\langle(a\otimes x)\xi,\xi\rangle=\tau_{A}(a)\tau_{Q}(x)$, and $\langle y\xi b,\xi\rangle=\tau_{Q}(y)\tau_{B}(b)$ for every $a\in A,~{}x,y\in Q,$ and $b\in B$.
@@
We prove in Proposition 3.5 that vNOE is indeed an equivalence relation. We should remark that, in the above definition, $\mathcal{H}$ can also be considered as an $A-B\,\overline{\otimes}\,Q^{\mathrm{op}}$-bimodule satisfying conditions analogous to the two mentioned in the definition. This essentially is the reason for the symmetry of vNOE, even though the definition seems asymmetric at first. To prove transitivity, inspired by [BV, Lemma 5.11], we establish an equivalent characterization of vNOE in Theorem 3.1, and show in Theorem 3.10 that $\Gamma\sim_{\mathrm{vNOE}}\Lambda$ if and only if $L\Gamma\sim_{\mathrm{vNOE}}L\Lambda$. Since $L(\Gamma*\Lambda)\cong L\Gamma*L\Lambda$, Theorem 1.3 follows from the following theorem, which we prove in Section 3. We should remark that the above definition, as stated, depends on the choice of the traces $\tau_{A}$ and $\tau_{B}$. Outside of the case of finite factors, it is not immediately clear whether the above definition is independent of the choice of the traces for general finite von Neumann algebras.
@@
If $(A_{i},\tau_{A_{i}}),(B_{i},\tau_{B_{i}}),~{}i=1,2$ are tracial von Neumann algebras such that $(A_{i},\tau_{A_{i}})\sim_{\mathrm{vNOE}}(B_{i},\tau_{B_{i}}),~{}i=1,2$, then, $(A_{1}*A_{2},\tau_{A_{1}}*\tau_{A_{2}})\sim_{\mathrm{vNOE}}(B_{1}*B_{2},\tau_{B_{1}}*\tau_{B_{2}})$.
Similar to free products, one also has that the group von Neumann algebra of a graph product of groups is isomorphic to the (von Neumann algebraic) graph product of the group von Neumann algebras, and hence Theorem 1.5 follows from the following theorem.
@@
Let $\mathcal{G}=(V,E)$ be a simple graph with at most countably infinite vertices. Let $(A,\tau_{A})$ and $(B,\tau_{B})$ be two graph products over $\mathcal{G}$, with tracial vertex von Neumann algebras $\{(A_{v},\tau_{A_{v}})\}_{v\in V}$ and $\{(B_{v},\tau_{B_{v}})\}_{v\in V}$, respectively. If $(A_{v},\tau_{A_{v}})\sim_{\mathrm{vNOE}}(B_{v},\tau_{B_{v}})$ for every $v\in V$, then $(A,\tau_{A})\sim_{\mathrm{vNOE}}(B,\tau_{B})$.
Remark 1.9.
Since graph product over a totally disconnected graph, i.e., a graph with no edges, gives free product, Theorem 1.7 follows from Theorem 1.8. In particular, Theorem 1.8 shows that Theorem 1.7 holds for free products of countably many tracial von Neumann algebras similar to the result obtained by Gaboriau in [Gab05, $\mathbf{P}_{\mathrm{ME}}{\bf 6^{*}}$]. Furthermore, Horbez and Huang [HH22, Proposition 4.2] proved the stability of OE under taking graph products over finite simple graphs, but our result holds for an arbitrary, countably infinite graph product. We include a proof of Theorem 1.7 for two reasons. Firstly, the notation is a little less involved compared to the proof of Theorem 1.8. Secondly, for the convenience of a reader who might be interested in the result but is not familiar with graph products.
In Proposition 3.13, we show that vNOE tracial von Neumann algebras are vNE in the sense of [IPR24]. We should remark that vNE does not imply vNOE in general.
@@ Acknowledgements.
The authors would like to express deep gratitude to Ben Hayes for his valuable insights, continuous support, and constant encouragement throughout this project. This work has benefitted profoundly from discussions with Ben Hayes on numerous occasions and from him encouraging us to look into [BV]. The project began with the authors’ meeting at the “NCGOA Spring Institute 2023” held at the Vanderbilt University, and we thank the organizers of the conference for the opportunity. A.Wu would also like to acknowledge the support from the NSF CAREER award DMS ##\##214473. Finally, we would like to thank the anonymous referee for helpful suggestions and pointing out an argument to us that helped in simplifying the proof of Theorem 1.8 and extending it to the case of an arbitrary countably infinite graph product.